A house has a length of 1200cm and width of 990cm.What is the area and perimeter of the house?

Answer by checkley77(12844)

P=2L+2W
P=2*1,200+2*990
P=2,400+1,980=4,380 CM. ANS.
AREA=L*W
AREA=1,200*990
AREA=1,188,000 CM.^2 ANS.